OpenStudy (anonymous):

solve the rational equation. check your answer. 2c/c - 4 - 2 = 4/c + 5

OpenStudy (e.mccormick):

I think you need some ( ) or the equation editor.... but to guess, did you mean: \[\frac{2c}{c - 4} - 2=\frac{4}{c + 5}\]

OpenStudy (anonymous):

Yes, that is what I meant.

OpenStudy (e.mccormick):

Have you tried using multiplication to cancel out the fractions?

OpenStudy (anonymous):

I know I need to cross multiply, but I don't know which numbers to multiply.

OpenStudy (e.mccormick):

Well, you can't just cross multiply because of that -2. But you can multiply through by each deniminaotr.

OpenStudy (e.mccormick):

I mean you have the distribution to deal with:\[\frac{2c}{c - 4} - 2=\frac{4}{c + 5}\implies \\ ( c + 5)\left(\frac{2c}{c - 4} - 2\right)=\left(\frac{4}{c + 5}\right)( c + 5)\implies \\ \frac{( c + 5)2c}{c - 4} - ( c + 5)2=4\]

OpenStudy (e.mccormick):

Once you do the same thing to \(c-4\) you will have it without the fractions. Then you can get it all to one side of the equation, combined like terms, and solve for 0s. However, if any answer you find would invalidate the original equation, it is not a useable answer! You have fractions, so if the divisor would be 0 because of the answer you find....

OpenStudy (anonymous):

I know my final answer should be -14, I just need to show my work

OpenStudy (e.mccormick):

Well, keep going the way I started that. You will have all the work.

OpenStudy (e.mccormick):

If there is something about what I did that you do not understand, what is it? That way I can explain it.

OpenStudy (anonymous):

I don't know what to do next

OpenStudy (e.mccormick):

You see how I got rid of one denominator, right?

OpenStudy (anonymous):

Not really, I'm terrible with math

OpenStudy (e.mccormick):

OK, lets go back to fraction and numbers for a bit. If I multiply \(\frac{1}{2}\) by 2, what happens to it?

OpenStudy (anonymous):

It becomes 1

OpenStudy (e.mccormick):

Yes. And there is a process to that. The 2 goes up on top, and multiplies the 1. \(\frac{1}{2}\cdot 2=\frac{1\cdot 2}{2}\) But anything times 1 is itself, so that simplifies: \(\frac{1\cdot 2}{2}=\frac{2}{2}\) Then, because there is a 2 on the top and the bottom, it cancels: \(\frac{2}{2}=\frac{1}{1}\) Which as you pointed out is just 1. Now, what if we put a letter in there for a variable? It mmakes it a little different, but it is the same set of steps.

OpenStudy (e.mccormick):

If I multiply \(\frac{3}{x}\) by x, what happens to it? The x goes up on top, and multiplies the 3. \(\frac{3}{x}\cdot x=\frac{3\cdot x}{x}\) Then, because there is an x on the top and the bottom, it cancels: \(\frac{3x}{x}=\frac{3}{1}\) Now, 3 over 1 is just 3. Not that different, right?

OpenStudy (anonymous):

yeah, it's just confusing because it's such a big problem. I never know what to multiply first

OpenStudy (e.mccormick):

No problem. It is getting there. I just want to show one last like thing this leads to.

OpenStudy (e.mccormick):

So now that you have seen this with a variable, what about a term, an expression, or what you might call a factor! What if I have \(\frac{4}{x + 5}\) and I multiply it by \(x + 5\)? \[\frac{4}{x + 5}\cdot ( x + 5)\implies \frac{4\cdot ( x + 5)}{x + 5}\implies \frac{4}{1}\implies 4\]Do you see how that worked?

OpenStudy (anonymous):

Yes

OpenStudy (e.mccormick):

Oops. Pasredin your problem wrong. LOL. Fixing that: Now that we have reviewed fractions a bit, lets see how that applies to your problem. So I looked at your problem:\[\frac{2c}{c - 4} - 2=\frac{4}{c + 5}\]And I thought to myself, "There are fractions. Only two ways to deal with fractions. Common denominators or eliminate the fractions." Personally, I like fractions. They are very useful. However, in this case you need to solve for c in the end. To solve for c, you must eliminate the fractions at some point. So that told me the best path to take is eliminating the fractions. To eliminate the \( c + 5\) on the right, I can do exactly what I did above! However, anything I multiply on the right, I MUST multiply on the left, so I have to multiply both sides by \(c + 5\).

OpenStudy (e.mccormick):

That was why my first step was:\[\frac{2c}{c - 4} - 2=\frac{4}{c + 5}\implies \\ ( c + 5)\left(\frac{2c}{c - 4} - 2\right)=\left(\frac{4}{c + 5}\right)( c + 5)\implies \\ \frac{( c + 5)2c}{c - 4} - ( c + 5)2=4 \]The second step is to do something similar to get rid of \(c-4\). Can you do that second step now?

OpenStudy (anonymous):

What do I need to do to get rid of the c - 4?

OpenStudy (e.mccormick):

Well, look at how I multiplied through the c+5. That is what you need to do, but with c-4.

OpenStudy (anonymous):

what do I multiply it by?

OpenStudy (e.mccormick):

Well, in every example with fractions above I got rid of the fraction by multiplying by what was on the bottom of the fraction. Does that explain how you choose what to multiply by?

OpenStudy (anonymous):

Not really. I would need to see it in the problem

OpenStudy (e.mccormick):

\[\frac{( c + 5)2c}{c - 4} - ( c + 5)2=4\]\(\uparrow\uparrow\uparrow\uparrow\uparrow\uparrow\uparrow\uparrow\uparrow\)There is is.

OpenStudy (anonymous):

so multiply it by itself on the left and it by (c+5) on the right?

OpenStudy (e.mccormick):

Well, there are three pieces of this equation. \[\text{One is:}\frac{( c + 5)2c}{c - 4} \text{ The second is}- ( c + 5)2 \text{ And there is the }4\]All three of these need to be multiplied by \(c-4\) to get rid of that fraction.

OpenStudy (anonymous):

I don't know how to distribute

OpenStudy (e.mccormick):

Well, did you see how I distributed when I did the part I did?

OpenStudy (anonymous):

I keep looking it over but it doesn't really make sense

OpenStudy (anonymous):

I still do not understand how to do it with the c-4

OpenStudy (e.mccormick):

Well, see how my pruple arrows go to three places? Same sort of distribution.

OpenStudy (anonymous):

I see where the arrows go but I don't really know how to use the information to do it with the c - 4

OpenStudy (e.mccormick):

Same basic procedure:

OpenStudy (anonymous):

I don't know how to multiply c- 4 if it were just 4 I could do it

OpenStudy (e.mccormick):

Hmm.... do you know how to do FOIL for things like (x+2)(x+2)?

OpenStudy (anonymous):

No

OpenStudy (e.mccormick):

OK. You saw the arrows, right? On the second distribute picture. That is the path the entire \((c-4)\) nees to follow.

OpenStudy (e.mccormick):

With the ( ) around it. They are important still.

OpenStudy (anonymous):

How is that relevant?

OpenStudy (e.mccormick):

Because they let you know that it still needs to be distributed after. ( ) are a way of saying something needs to be multiplied. For example, when I did this: \[\frac{( c + 5)2c}{c - 4} - ( c + 5)2=4\] The \((c+5)\) still needs to be multiplied by a few things there.

OpenStudy (anonymous):

This is exactly where i am at. Could you please walk me through the rest of this problem so I can try to understand how to finish the rest my my assignment?

OpenStudy (e.mccormick):

That is what I have been trying to do. I was going to take care of the divisors first, then finish multiplying things out. So, can you use the arrows in the second picture to distribute the \((c-4)\)? The arrows show you where it needs to go.